OpenStudy (anonymous):

Chico walks 3/5 mile to the library. Nikki walks 1/4 the distance that Chico walks to the library. How far does Nikki walk?

OpenStudy (anonymous):

Just multiply the fractions.\[\frac{ 3 }{ 5 }*\frac{ 1 }{ 4} \to \frac{ 3*1 }{ 5*4 } = \frac{ 3 }{ 20 }\] Since we can't simplify it any further, the answer would simply be 3/20. Any questions as to how I got it?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Ok, then good